# ALiiCE: Evaluating Positional Fine-grained Citation Generation

### v1

@@ Abstract
Large Language Models (LLMs) can enhance the credibility and verifiability by generating text with citations. However, existing tasks and evaluation methods are predominantly limited to sentence-level statement, neglecting the significance of positional fine-grained citations that can appear anywhere within sentences. To facilitate further exploration of the fine-grained citation generation, we propose ALiiCE, the first automatic evaluation framework for this task. Our framework first parses the sentence claim into atomic claims via dependency analysis and then calculates citation quality at the atomic claim level. ALiiCE introduces three novel metrics for positional fined-grained citation quality assessment, including positional fine-grained citation recall and precision, and coefficient of variation of citation positions. We evaluate the positional fine-grained citation generation performance of several LLMs on two long-form QA datasets. Our experiments and analyses demonstrate the effectiveness and reasonableness of ALiiCE. The results also indicate that existing LLMs still struggle to provide positional fine-grained citations. [^1] Our code is available at https://github.com/ylXuu/ALiiCE.
@@ 1 Introduction
Large Language Models (LLMs; Brown et al., 2020) can improve performance in several NLP tasks by incorporating external knowledge  (Lewis et al., 2020). In order to improve LLMs’ credibility,  Gao et al. (2023b); Liu et al. (2023) propose a new paradigm for long-form QA, which LLMs are required to provide citations to the retrieved passages for the statements they generate. Since then, many studies  (Ye et al., 2024; Huang et al., 2024; Slobodkin et al., 2024) have focused on how to enhance LLMs’ citation generation capabilities.
However, existing methods and evaluation metrics on citation generation are predominantly limited to sentence-level statements.  Malaviya et al. (2024) suggest that a sentence might not be the smallest unit capable of representing an atomic claim, potentially leading to inaccurate evaluations. As illustrated in Figure 1, the response A1 actually contains two different claims, but the sentence-level citation treats the entire sentence as one claim. Additionally,  Liu et al. (2023) highlight that the generated text scope of a single in-line citation is often ambiguous. Citations of A1 in Figure 1 is ambiguous, because the citation marks at the end of A1 do not clearly indicate whether they support both claims or only the last claim.
@@
In fact, in many long-form contexts, particularly in professional fields such as academic writing  (Funkquist et al., 2023), citation marks often appear in the middle of a sentence rather than always at the end, as response A2 shown in Figure 1. Compared with sentence-level citation, the advantages of this fine-grained generation are: 1) clearer indication of the text scope associated with each citation mark, and 2) better user-friendliness, allowing users to locate more specific content to check. We refer to this improved generation task as Positional Fine-grained Citation Text Generation.
@@
Despite the importance of this task, an effective evaluation method has yet to be developed. Sentence-level metrics simply merge citations from different positions, treating the entire sentence as a single claim  (Gao et al., 2023b; Yue et al., 2023). In this case, when using Natural Language Inference (NLI; Honovich et al., 2022) to judge whether the claim is supported by its evidences, sentence-level metrics can easily result in the issue of excessively long NLI contexts when multiple atomic claims occur simultaneously within the sentence. Furthermore, if there is an overlap between evidence of different atomic claims, sentence-level judgments can also become unreasonable, for correct citations might be mistakenly excluded.
@@
To effectively assess fine-grained citations at atomic claim level, we propose ALiiCE, representing Automatic LLM’s Positional Fine-grained Citation Evaluation. Our method first employs a Dependency Tree based approach to parse the atomic claim corresponding to each citation in the response. For instance, the two claims of sentence A2 in Figure 1 are parsed as "Cups can be made of glass" and "Cups can be made of plastic". Further, our method incorporates three new metrics on citation quality, including positional fine-grained citation recall and precision, as well as coefficient of variation of citation positions for assessing the dispersion of citation placements within a sentence.
We conduct experiments to evaluate the performance of existing models on positional fine-grained citation generation. We employ two long-form QA datasets, ASQA  (Stelmakh et al., 2022) and ELI5  (Fan et al., 2019) to evaluate LLMs including GPT-3.5, GPT-4 and LLaMA-3-8B.
Through the analyses on experimental results and cases, we demonstrate that ALiiCE is an effective and reasonable evaluation method for positional fine-grained citation generation. We observe that existing LLMs face challenges in generating positional fine-grained citations. These challenges encompass producing a limited number of fine-grained citations and struggling to generate accurate ones. In addition, we find that the latest open-source LLMs narrow the gap in citation generation with closed-source LLMs. We hope that our work can inspire more research into positional fine-grained citation text generation.
In summary, our main contributions include:
  • We propose a dedicated evaluation method for positional fine-grained citation generation.
  • We analyze and verify the shortcomings of sentence-level evaluation in positional fine-grained citation text generation, as well as the effectiveness of our method.
  • We evaluate the performance of existing LLMs on positional fine-grained citation text generation on long-form QA datasets.
@@ 2.1 Citation Generation in Long-form QA
Long-form QA is a type of Question-Answering task, where the answer to a question is detailed, comprehensive, and typically longer than brief answers. Unlike short-form QA, which typically provide binary, entity-level or short sentence answers, long-form QA generates more elaborate responses that include explanations, context, and additional relevant information (Krishna et al., 2021).
Citation generation involves producing citation marks (namely, document IDs) while generating text, indicating the source documents on which the text is based  (Funkquist et al., 2023). In our work, we focus on citation generation for long-form QA. Unlike traditional task, positional fine-grained citation generation allow citation marks to appear at any position within the sentence.
@@ 2.2 Task Definition
Formally, given a query $q$ and a set $\mathcal{D}$ of retrieved passages based on $q$, the generator $\mathcal{M}$ is required to generate a response $\mathcal{R}$ consisting of $n$ sentences $s_{1},\dots,s_{n}$. We assume that the $j$-th item in $i$-th sentence $s_{i}$ is $s_{i,j}$, which has two situations:
$s_{i,j}=\begin{cases}\mathcal{C}_{i,j},&\text{if $s_{i,j}$ is a group of citation marks}\\ x_{i,j},&\text{if $s_{i,j}$ is a word}\end{cases}$ (1)
where $\mathcal{C}_{i,j}=\left\{c_{i,j,1},c_{i,j,2},\dots\right\},c_{i,j,k}\in\mathcal{D}$. If a $s_{i,j}$ is a group of citation marks, then it has its corresponding claim generated by $\mathcal{M}$ based on $\mathcal{C}_{i,j}$, denoted as $\mathcal{A}_{i,j}$. Obviously, $\mathcal{A}_{i,j}$ is constructed from $\left\{x_{i,j}\right\}$, as a part of sentence $s_{i}$.
Take A2 in Figure 1 as an example, "plastic" is a word, and "[2][3]" is a group of citation marks with the claim "Cups can be made of plastic".
@@ 3 ALiiCE: Automatic LLMs’ positional fine-grained Citation Evaluation
In this section, we give a detailed description of ALiiCE. First, we introduce how we construct the atomic claim parsing pipeline based on dependency trees and Lowest Common Ancestor (LCA). Then, we present our novel metrics for the evaluation of positional fine-grained citation quality.
@@ 3.1 Dependency Tree
Dependency tree is a hierarchical representation of the grammatical structure of a sentence, showing how words rely on each other  (Culotta and Sorensen, 2004). Compared with the hierarchical syntax tree based on operators, the dependency tree is more concise, making it easier to extract the relationship between sentence components. And in a dependency tree, a subtree can represent a phrase or clause that depends on its root. Thus, dependency tree is highly suitable for extracting the claims associated with the citation marks.
In ALiiCE, we employ dependency trees to describe sentences in $\mathcal{R}$ for subsequent parsing stage. In practice, for every $\mathcal{C}_{i,j}$, we match a node of word $x_{i,j}$ in the tree of sentence $s_{i}$, so that every node in dependency tree is a word.
3.2 Lowest Common Ancestor
Lowest Common Ancestor (LCA) is the lowest (i.e. deepest) node of two different nodes possessing both of them as descendants in a tree. When there are multiple citations in different locations within a sentence, their respective claims need to be parsed. To parse a claim of $\mathcal{C}_{i,j}$, we need to exclude irrelevant content from other claims, as different claims may share identical sentence components. In the dependency tree, for two distinct nodes, their dependency traces back to their LCA node. The subtrees of the LCA node’s children, containing two word nodes matched by two citations, represent two components that depend on the LCA node. Based on the relationship between two subtrees, we can delete or modify one of them, thereby achieving claim extraction of the citation of the other subtree.
@@ 3.3 Parsing Pipeline
For a sentence $s_{i}$ in response $\mathcal{R}$, we extract groups of citation marks $\left\{\mathcal{C}_{i,j}\right\}$ from difference positions. Then we do text cleaning on $s_{i}$ to obtain raw sentence ${s_{i}}^{\prime}$, involving removing citation marks and other punctuation. ${s_{i}}^{\prime}$ is used to construct dependency tree $T$. Next, for each $\mathcal{C}_{i,j}$, we match a word node, which we denote as a citation node, for simplicity. The principle of matching nodes is to select the $x_{i,j}$ closest to $\mathcal{C}_{i,j}$ in $s_{i}$, giving priority to the one before $\mathcal{C}_{i,j}$. Thus we modify the dependency tree based on the citation nodes.
@@
After iteration, we obtain a modified dependency tree. We convert words in the modified tree to text following the order in original sentence, getting the claim of citations corresponding to node $i$. We provide additional details and examples of our algorithm in Appendix A and B, respectively.
@@
Input: A sentence $s$ with inline citation marks
@@
3.4 Metrics For Fine-grained Citation
In this section, we display three new metrics for positional fine-grained citation quality in ALiiCE.
@@ 3.4.1 Positional Fine-grained Citation Recall
For each $\mathcal{C}_{i,j}$ and its corresponding $\mathcal{A}_{i,j}$, if the concatenation of passages in $\mathcal{C}_{i,j}$ can entail $\mathcal{A}_{i,j}$, then the citation recall of $\mathcal{C}_{i,j}$ is 1, otherwise it is 0. The judgement of entailment can be formulated as:
@@ 3.4.2 Positional Fine-grained Citation Precision
Following  (Gao et al., 2023b), we calculate citation precision to evaluate whether every citation is necessary. This metric checks for redundant citations to improve readability and verifiability.
@@
We only calculate citation precision when the citation recall of $\mathcal{C}_{i,j}$ is 1. Specifically, for each $c_{i,j,k}$ in $\mathcal{C}_{i,j}$, if $c_{i,j,k}$ can not entail $\mathcal{A}_{i,j}$ alone while the concatenation of passages in $\mathcal{C}_{i,j}\setminus c_{i,j,k}$ can, it is indicated that $c_{i,j,k}$ is a redundant citation and the precision score of $c_{i,j,k}$ is 0, otherwise the precision score of $c_{i,j,k}$ is 1. Then we calculate the mean of the precision scores from each $c_{i,j,k}$ as the precision score of $\mathcal{C}_{i,j}$. If the citation recall of $\mathcal{C}_{i,j}$ is 0, then its citation precision is 0, as well.
@@
Model ($k$-psg-form) ALiiCE ALCE Fluency Correct. CPCV Length
Rec. Prec. Rec. Prec.
GPT-3.5 (5-psg) 78.4 (0.5) 74.4 (0.4) 80.0 (0.3) 72.9 (0.4) 86.1 (2.9) 51.1 (0.3) 0.10 (-) 50.5 (37.3)
GPT-3.5 (5-psg-summ) 76.9 (0.4) 71.6 (0.9) 77.3 (0.3) 71.5 (0.6) 75.4 (2.3) 49.3 (0.3) 0.13 (-) 40.2 (33.2)
GPT-3.5 (5-psg-snip) 74.4 (0.7) 69.4 (0.3) 74.8 (0.2) 68.2 (0.4) 73.1 (3.7) 48.0 (0.6) 0.13 (-) 36.0 (29.7)
GPT-3.5 (10-psg) 77.7 (1.2) 75.9 (0.8) 79.8 (1.5) 73.8 (1.4) 84.6 (7.1) 44.1 (0.3) 0.15 (-) 63.4 (52.6)
GPT-4 (5-psg) 76.8 (1.2) 68.2 (1.1) 78.5 (0.6) 67.0 (0.6) 52.2 (9.5) 47.0 (0.4) 0.15 (-) 28.1 (20.8)
LLaMA-3-8B (5-psg) 64.8 (1.0) 61.4 (1.4) 65.8 (1.9) 60.3 (1.2) 84.2 (5.0) 50.9 (0.3) 0.44 (-) 64.0 (53.1)
LLaMA-3-8B (10-psg) 61.8 (1.3) 62.5 (0.5) 62.0 (1.3) 60.1 (1.5) 88.8 (9.6) 41.7 (1.4) 0.45 (-) 73.2 (64.9)
@@
Table 1: Results on ASQA. The $k$-psg indicates using top-$k$ relevant documents for response generation. Document formats include summary (summ), snippet (snip), and default original text. The correctness here refers to the exact match recall. The value in bracket represents the population standard deviation.

### v2

@@ Abstract
Large Language Model (LLM) can enhance its credibility and verifiability by generating text with citations. However, existing research on citation generation is predominantly limited to sentence-level statements, neglecting the significance of positional fine-grained citations that can appear anywhere within sentences. To facilitate further exploration of the positional fine-grained citation generation, we propose ALiiCE, the first automatic evaluation framework for this task. Our method employs a dependency tree based approach to parse the sentence-level claim into atomic claims. Then ALiiCE evaluates citation quality using three metrics, including positional fine-grained citation recall, precision, and coefficient of variation of citation positions. We evaluate the positional fine-grained citation generation performance of several LLMs on long-form QA datasets. Our experiments and analyses demonstrate the effectiveness and reasonableness of ALiiCE. We offer our insights into the current advancements and future directions for the positional fine-grained citation generation task. [^1] Our code is available at https://github.com/ylXuu/ALiiCE.
@@ 1 Introduction
Large Language Models (LLMs; Brown et al., 2020) can improve performance in several NLP tasks by incorporating external knowledge  (Lewis et al., 2020). In order to improve LLMs’ credibility,  Gao et al. (2023b); Liu et al. (2023) propose a new paradigm for long-form QA, in which LLMs are required to provide citations to the retrieved passages for the statements they generate. Since then, many studies  (Ye et al., 2024; Huang et al., 2024a; Slobodkin et al., 2024) have focused on how to enhance LLMs’ citation generation capabilities.
However, existing research on citation generation is predominantly limited to sentence-level statements.  Malaviya et al. (2024) suggest that a sentence might not be the smallest unit capable of representing an atomic claim, potentially leading to inaccurate evaluations. As shown in Figure 1, response A1 actually contains two different claims, but the sentence-level citation treats the entire sentence as one claim. Additionally,  Liu et al. (2023) highlight that the generated text scope of a single in-line citation is often ambiguous. Citations of A1 in Figure 1 is ambiguous, because the citation marks at the end of A1 do not clearly indicate whether they support both claims or only the last claim.
@@
In fact, in many long-form contexts, particularly in professional fields such as academic writing  (Funkquist et al., 2023), citation marks often appear in the middle of a sentence rather than always at the end, as response A2 illustrated in Figure 1. Compared with sentence-level citation, the advantages of this fine-grained generation are: 1) clearer indication of the text scope associated with each citation mark, and 2) better user-friendliness, allowing users to locate more specific content to check. We refer to this improved generation task as Positional Fine-grained Citation Text Generation.
@@
Despite the importance of this task, an effective evaluation method has yet to be developed. Some studies directly apply sentence-level metrics to fine-grained citations (Huang et al., 2024b), but this can affect the accuracy of evaluation. First, sentence-level metrics simply merge evidences of multiple atomic claims  (Gao et al., 2023b). When using Natural Language Inference (NLI; Honovich et al., 2022) to judge entailment, merged evidences can easily result in the issue of excessively long NLI contexts. Second, if there is an overlap between evidence of different atomic claims, sentence-level judgments can also become unreasonable, for correct citations might be mistakenly excluded. Thus, it is essential to design an evaluation method specifically tailored for positional fine-grained citations.
@@
We propose a new evaluation method, ALiiCE, Automatic LLM’s Positional Fine-grained Citation Evaluation. Our method first employs a Dependency Tree based approach to parse atomic claim of each citation in the response. For instance, the two claims of A2 in Figure 1 are parsed as "Cups can be made of glass" and "Cups can be made of plastic". Further, our method incorporates three metrics for evaluating citation quality, including citation recall and precision at the level of atomic claims, as well as the Coefficient of Variation of Citation Positions (CVCP), which measures the dispersion of citation positions within a sentence.
In our experiment, we employ two long-form QA datasets, ASQA  (Stelmakh et al., 2022) and ELI5  (Fan et al., 2019) to evaluate outputs of LLMs including GPT-3.5, GPT-4 and LLaMA-3-8B. We observe that existing LLMs generate a limited number of positional fine-grained citations. We compare the citation quality of LLMs’ output in sentence-level metrics with ALiiCE to demonstrate the necessity of evaluation method for positional fine-grained citations. We also conduct error analyses to assess the impact of parsing errors. Additionally, we conduct human evaluation to verify the consistency between ALiiCE and human judgment.
Our main contributions include: 1) we propose the first dedicated evaluation method for positional fine-grained citation generation and we prove its effectiveness through experiments; 2) we evaluate the performance of several existing LLMs on positional fine-grained citation generation in long-form QA datasets; 3) we offer our insights on the study of positional fine-grained citation generation. We hope that our work can inspire more research into positional fine-grained citation text generation.
@@ 2.1 Citation Generation In Long-form QA
Unlike short-form QA, which typically provide binary, entity-level or short sentence answers, long-form QA generates longer responses including explanations, context, and additional relevant information (Krishna et al., 2021). Citation generation involves producing citation marks (namely, passage IDs) while generating text, indicating the source passages on which the text is based  (Funkquist et al., 2023). In our work, we focus on positional fine-grained citation generation for long-form QA. Unlike traditional task, it allow citation marks to appear at any position within the sentence.
@@ 2.2 Task Definition
Formally, given a query $q$ and a set $\mathcal{D}$ of retrieved passages based on $q$, the generator $\mathcal{M}$ is required to generate a long-form response $\mathcal{R}$ containing citations. Specifically, $\mathcal{R}$ is composed of several sentences, with each sentence containing words and in-line citation markers. We assume that the $k$-th sentence $s_{k}$ has a length of $l$ and can be represented as $x_{1},x_{2},\ldots,x_{l}$, where $x_{i}$ represents the $i$-th minimal semantic unit in $s_{k}$.
In this paper, the minimal semantic unit can be either a word (including punctuation) or a group of citation marks. If $x_{i}$ is the latter, we denote it as $\mathcal{C}_{i}=\left\{c_{i,1},c_{i,2},\ldots\right\}$, where $c_{i,j}$ is a citation mark of a passage in $\mathcal{D}$. And $C_{i}$ corresponds to an atomic claim parsed from its sentence, marked as $\mathcal{A}_{i}$. Take A2 in Figure 1 as an example, "plastic" is a word, and "[2][3]" is a group of citation marks with its atomic claim "Cups can be made of plastic".
@@ 3 ALiiCE: Automatic LLMs’ Positional Fine-grained Citation Evaluation
In this section, we give a detailed description of ALiiCE. First, we introduce how we construct the atomic claim parsing pipeline based on dependency trees. Then, we present three metrics for the evaluation of positional fine-grained citation quality.
@@ 3.1 Dependency Tree
Dependency tree is a hierarchical representation of the grammatical structure of a sentence, showing how words rely on each other  (Culotta and Sorensen, 2004), and is more concise compared with the hierarchical syntax tree based on operators. In a dependency tree, a subtree can represent a phrase or clause that depends on its root, which is highly suitable for atomic claims extraction. Thus in ALiiCE, we employ dependency trees to represent sentences in $\mathcal{R}$ for subsequent parsing stage.
For simplicity, we assume that the nodes in the dependency tree are all words. To extract atomic claim based on the position of the $\mathcal{C}_{i}$ in the original sentence, we find a matching node in the tree for each $\mathcal{C}_{i}$, as shown in the lower left part of Figure 2. In this paper, we refer to the node with citation marks attached as the citation node.
When handling multiple citations in different positions within a sentence, their respective claims need to be parsed. To parse the claim of $\mathcal{C}_{i}$, we need to exclude irrelevant content from other claims, as different claims may share identical sentence components. Thus, we consider the Lowest Common Ancestor (LCA), which is the deepest node of two different nodes possessing both of them as descendants in a tree. For two distinct citation nodes, we can modify the dependency tree to obtain atomic claims based on the relative positions with respect to their LCA node (see Section 3.2).
@@ 3.2 Parsing Pipeline
For sentence $s_{k}$ in response $\mathcal{R}$, we extract groups of citation marks $\left\{\mathcal{C}_{i}\right\}$ from difference positions. Then we do text cleaning on $s_{k}$ to obtain raw sentence ${s_{k}}^{\prime}$, involving removing citation marks and other punctuation. ${s_{k}}^{\prime}$ is used to construct dependency tree $T$. Next, we match citation node for each $\mathcal{C}_{i}$. The principle of matching node is to select word closest to $\mathcal{C}_{i}$ in $s_{k}$, giving priority to the one before $\mathcal{C}_{i}$. Then we modify the dependency tree based on the citation nodes.
@@
After iteration, we obtain a modified dependency tree. We convert words in the modified tree to text following the order in original sentence, getting the claim of citations corresponding to node $i$. We provide additional details and running examples of our algorithm in Appendix C and D, respectively.
@@
Input: A sentence $s$ with in-line citation marks
@@
3.3 Metrics For Citation Quality
In this section, we display the three metrics for positional fine-grained citation quality in ALiiCE.
@@ 3.3.1 Positional Fine-grained Citation Recall
For each $\mathcal{C}_{i}$ and its corresponding $\mathcal{A}_{i}$, if the concatenation of passages in $\mathcal{C}_{i}$ can entail $\mathcal{A}_{i}$, then the citation recall is 1, otherwise it is 0. The judgement of entailment can be formulated as:
@@ 3.3.2 Positional Fine-grained Citation Precision
Following  Gao et al. (2023b), we calculate citation precision to evaluate whether every citation is necessary. This metric checks for redundant citations to improve readability and verifiability.
@@
We compute citation precision only when the citation recall of $\mathcal{C}_{i}$ is 1; otherwise, the citation precision is set to 0. Specifically, for each $c_{i,j}$ in $\mathcal{C}_{i}$, if $c_{i,j}$ can not entail $\mathcal{A}_{i}$ alone while the concatenation of passages in $\mathcal{C}_{i}\setminus c_{i,j}$ can, it is indicated that $c_{i,j}$ is a redundant citation and the precision score of $c_{i,j}$ is 0, otherwise the precision score of $c_{i,j}$ is 1. Finally we calculate the mean of the precision scores from each $c_{i,j}$ as the precision score of $\mathcal{C}_{i}$.
3.3.3 Coefficient of Variation of Citation Positions
Positional fine-grained citation generation allows citation marks to appear in multiple positions within a sentence (e.g., in the middle, at the end). Consequently, to some extent, the dispersion of citation marker positions can reflect the LLMs’ ability to generate positional fine-grained citations. For example, in Figure 1, A2 has a greater dispersion of citation marker positions than A1. To quantify the degree of dispersion, we propose CVCP (Coefficient of Variation of Citation Positions).
For response $\mathcal{R}$, we first calculate the indices of citation marks’ positions for every sentence. For sentence $s_{k}$, which has a length of $l$ and can be represented as $x_{1},\ldots,x_{l}$, we extract the subscripts corresponding to the citation marks as the indices, denoted by $p_{1},\ldots,p_{t}$, where $t$ is the number of citation marks. We normalize the indices to eliminate the interference of sentence length as follows:
$p_{i}\leftarrow\frac{p_{i}}{\left|s_{k}\right|}$ (2)
Then we compute standard deviation for $s_{k}$ as:
[…]
where $\mu_{k}=\frac{1}{t}\sum^{t}_{j=1}p_{j}$, which represents the mean of normalized indices. Assuming $s_{k}$ has $n$ sentences, the CVCP of $\mathcal{R}$ is as follows:
$CV_{CP}\left(\mathcal{R}\right)=\frac{1}{n}\sum^{n}_{k=1}\frac{\sigma\left(s_{k}\right)}{\mu_{k}}$ (4)
When the positions of the citation markers in the sentence are more dispersed, the CVCP can be higher. Conversely, if all citation markers appearing at the end of the sentence, the CVCP can be very low (i.e., 0). Thus, CVCP encourages LLMs to generate more positional fine-grained citations.
@@ 3.3.3 Coefficient of Variation of Citation Positions
Dataset Model ($k$-psg-form) ALiiCE CVCP Fluency Correct. Length
Rec. Prec. F1.
ASQA GPT-3.5 (5-psg) 78.4 (0.5) 74.4 (0.4) 76.3 (-) 0.10 (-) 86.1 (2.9) 51.1 (0.3) 50.5 (37.3)
GPT-3.5 (5-psg-summ) 76.9 (0.4) 71.6 (0.9) 74.2 (-) 0.13 (-) 75.4 (2.3) 49.3 (0.3) 40.2 (33.2)
GPT-3.5 (5-psg-snip) 74.4 (0.7) 69.4 (0.3) 71.8 (-) 0.13 (-) 73.1 (3.7) 48.0 (0.6) 36.0 (29.7)
GPT-3.5 (10-psg) 77.7 (1.2) 75.9 (0.8) 76.8 (-) 0.15 (-) 84.6 (7.1) 44.1 (0.3) 63.4 (52.6)
GPT-4 (5-psg) 76.8 (1.2) 68.2 (1.1) 72.2 (-) 0.15 (-) 52.2 (9.5) 47.0 (0.4) 28.1 (20.8)
LLaMA-3-8B (5-psg) 64.8 (1.0) 61.4 (1.4) 63.1 (-) 0.44 (-) 84.2 (5.0) 50.9 (0.3) 64.0 (53.1)
LLaMA-3-8B (10-psg) 61.8 (1.3) 62.5 (0.5) 62.1 (-) 0.45 (-) 88.8 (9.6) 41.7 (1.4) 73.2 (64.9)
ELI5 GPT-3.5 (5-psg) 61.0 (0.5) 58.6 (2.2) 59.8 (-) 0.10 (-) 21.8 (0.6) 20.8 (0.3) 131.7 (46.2)
GPT-3.5 (5-psg-summ) 53.9 (2.6) 52.0 (1.1) 52.9 (-) 0.15 (-) 21.3 (5.1) 20.8 (1.1) 111.3 (46.5)
GPT-3.5 (5-psg-snip) 53.4 (1.3) 50.9 (1.1) 52.1 (-) 0.13 (-) 34.9 (7.3) 20.8 (0.4) 106.7 (47.9)
GPT-3.5 (10-psg) 58.1 (2.4) 56.8 (2.0) 57.4 (-) 0.12 (-) 18.5 (4.7) 19.7 (0.7) 155.9 (57.4)
GPT-4 (5-psg) 55.1 (0.5) 54.0 (3.0) 54.5 (-) 0.15 (-) 20.4 (7.2) 21.3 (0.9) 102.2 (59.7)
LLaMA-3-8B (5-psg) 45.9 (0.3) 47.1 (0.7) 46.5 (-) 0.53 (-) 36.2 (1.0) 20.5 (0.9) 203.9 (71.4)
LLaMA-3-8B (10-psg) 42.8 (0.8) 44.2 (0.9) 43.5 (-) 0.61 (-) 32.5 (6.2) 19.5 (0.7) 224.2 (77.7)
@@
Table 1: Results on ASQA and ELI5. The $k$-psg indicates using top-$k$ relevant documents for response generation. Document formats include summary (summ), snippet (snip), and default original text. The correctness refers to the exact match recall for ASQA and ROUGE-L for ELI5. The value in bracket represents the standard deviation.
4 Experimental Setup
In this section, we describe the datasets and implementation details of our experiments. Additional details are provided in Appendix A.
Datasets
We utilize two popular datasets for the task of long-form QA, including: 1) ASQA, which is an open-domain long-form QA dataset for ambiguous factoid queries, collected from AmbigQA  (Min et al., 2020); 2) ELI5, which is a dataset for complex QA with paragraph-length responses, collected from subreddit "Explain Like I’m Five". The queries of these two datasets are well suited for retrieval-augmented generation, thus more conducive for evaluating fine-grained citation generation. Following Gao et al. (2023b), we use the Generalizable T5-based dense Retriever  (GTR; Ni et al., 2022) to retrieve relevant passages for queries from Wikipedia corpus snapshot dated 2018-12-20.
Implementation
We utilize SpaCy [^2] https://spacy.io/ to construct dependency trees for sentences, which is a useful and efficient python toolkit for many NLP tasks. We use TRUE [^3] https://huggingface.co/google/t5_xxl_true_nli_mixture, a fine-tuned T5-11B  (Raffel et al., 2020) model as the NLI model for the judgement of entailment in citation quality.
Models
For closed-source LLMs, we evaluate gpt-4-turbo-2024-04-09 and gpt-3.5-turbo-0125  (OpenAI, 2022; OpenAI et al., 2024). For open-source LLMs, we evaluate LLaMA-3-8B  (AI@Meta, 2024). In addition, we incorporate variables such as the number of retrieved passages and the passage form used in generation (truncated original text, summary, or snippet) into the model setting. The prompts are provided in Appendix E.
Evaluation Metrics
In addition to the three metrics of citation quality introduced at Section 3.3, we utilize three common metrics in long-form QA, including: 1) correctness, which checks whether $\mathcal{R}$ answers the query $q$ accurately; 2) fluency, which evaluates whether $\mathcal{R}$ is coherent; and 3) length, which is the average length of $\mathcal{R}$. For comparison, we also use ALCE (Gao et al., 2023b) as the sentence-level evaluation to assess citation quality.
5 Main Results
In this section, we present our key observations on the experiment results, and then provide our case study to prove the necessity of developing method for positional fine-grained citation evaluation.
5.1 Overall Performances
The result of our experiment is presented in Table 1. We obtain some key observations as follows:
Citation quality
In ASQA, GPT-3.5 (10-psg) achieves the best performance in citation recall and precision, while in ELI5, the top performer is GPT-3.5 (5-psg). Overall, these two models exhibit outstanding performance of citation quality across both datasets. Moreover, simpler passage formats, such as summary and snippet, do not yield performance improvements. Through CVCP, we observe that most models generate a limited number of positional fine-grained outputs. LLaMA-3-8B is able to generate more fine-grained samples than GPT-3.5 and GPT-4, among which LLaMA-3-8B (10-psg) achieves the highest CVCP in both datasets.
Other metrics
The difference in fluency between the models is not significant; however, it is evident that the model outputs for ELI5 are much less fluent compared to ASQA. This discrepancy is also observed in terms of correctness. Regarding length, the model outputs for ELI5 are substantially longer than those for ASQA, and the output length of LLaMA-3 is longer than that of GPT-3.5 and GPT-4. As for the difference between the two datasets, we believe that this may be due to the more difficult queries and the more complex knowledge contained in the passages in ELI5.
Figure 3: Evaluation process of citation quality by ALCE and ALiiCE on two examples from ASQA. The answers are generated by GPT-3.5 (5-psg).
5.2 Case Study
In this section, we compare the evaluations of ALCE and ALiiCE on two instances from ASQA, and analyze the shortcomings and insufficiency of sentence-level metrics on positional fine-grained citation evaluation. Our objective is to demonstrate the necessity of designing a dedicated citation evaluation method with atomic claim parsing.
Long-context issue
Citation precision issue
If there is an overlap between different evidences, it is potential for the NLI model to misjudge multiple atomic claims simultaneously. Taking the Case 2 in Figure 3 as an example, citation "[3]" contains evidences supporting both atomic claim 1 and 2. According to ALCE’s citation precision, citation "[3]" alone can support the entire sentence-level claim, whereas citation "[4]" cannot, as it only supports atomic claim 2. Consequently, citation "[4]" is considered redundant, despite being or even though it is actually a reasonable citation. In ALiiCE, we evaluate based on atomic claims, ensuring that the assessment is not influenced by evidences from other claims.
Dataset Num of Claims Num of same NLI
ASQA 1935 1930
ELI5 3923 3891
5.3 Error Analyses
Grammatical error
Parsing error
Therefore, we conduct further experiment to test the potential impact of parsing errors on NLI. We firstly collect all the atomic claims from two datasets. Next, we utilize GPT-3.5 to refine each claim based on its original sentence (the prompt is provided in Appendix E). And then we employ the NLI model to assess the entailment before and after the claim refinement. As indicated in Table 2, the result show that the proportion of claims with inconsistent NLI results is less than 1% across both datasets. Therefore, the parsing error is unlikely to have a significant impact on the evaluation.